OpenStudy (anonymous):

1.05] Use the distributive property to simplify the expression below: −2(5d − 1) + 4d −6d − 2 −10d + 2 14d + 2 −6d + 2

OpenStudy (anonymous):

did you try this?

OpenStudy (ajprincess):

Distributive property a(b+c)=a*b+a*c=ab+ac Can u do nw?